# Parsimonious Compactly Supported Covariance Models in the Gauss Hypergeometric Class: Identifiability, Reparameterizations, and Asymptotic Properties

### v1

@@ 5 Some special cases and bridges to the Matérn model
The computation of the $\mathcal{H}_{\kappa,\mu,a,d}$ correlation function requires an efficient evaluation of the Gauss hypergeometric function, which is feasible (Johansson,, 2019) but could be cumbersome to (spatial) statisticians or machine learning practitioners used to handle closed-form parametric correlation models.
However, similarly to the $\mathcal{MT}$ and $\mathcal{GW}$ models, in some special cases, the computation can be considerably simplified. The following theorem considers the case when $\kappa$ is a non-negative integer, $d=1$ and $\mu\geq 1$.
@@
If $\kappa=k$ is a non-negative integer, $d=1$ and $\mu\geq 1$, then:
$\begin{split}\mathcal{H}_{k,\mu,a,1}(x)&=\frac{\sqrt{\pi}\Gamma\left(\frac{\mu}{2}+2k+1\right)}{\Gamma\left(k+0.5\right)}\left(1-\frac{x}{a}\right)^{\mu+2k}\left(\frac{x}{a}\right)^{k}\\ &\times\sum\limits^{k}_{n=0}\frac{(n+k)!}{2^{2n}n!\Gamma(n+\frac{\mu}{2}+k+1)(k-n)!}\left(\frac{(1-\frac{x}{a})^{2}}{\frac{x}{a}}\right)^{n}.\end{split}$
Unfortunately, general closed forms for any $d>1$, $\mu\geq 1$ and $\kappa$ a non-negative integer, are difficult to obtain. Some special cases can, however, be obtained using closed forms of the Gauss hypergeometric function. For instance, the following theorem provides a closed form expression when $d$ is an odd integer, $\mu\geq 1$ and $\kappa=0$.
@@
Let $d$ an odd integer and $\mu\geq 1$. If $\kappa=0$ then
$\mathcal{H}_{0,\mu,a,d}(x)=C\left(1-\frac{x}{a}\right)_{+}^{\mu+\frac{d-1}{2}}\sum\limits^{\frac{d-1}{2}}_{n=0}\frac{\left(\frac{1-d}{2}\right)_{n}\left(\frac{d-1}{2}+1\right)_{n}\left(1-\frac{x}{a}\right)^{n}_{+}}{2^{n}n!\Gamma\left(n+\frac{d-1}{2}+\mu+1\right)},$
with $C=\frac{2^{\mu+\frac{d-1}{2}}\Gamma\left(\frac{\mu+1}{2}\right)\Gamma\left(\frac{\mu+d+1}{2}\right)}{\sqrt{\pi}}$.
Additional closed-form expressions, including the spherical models, for some fixed $\mu$ and $\kappa$ are given in Appendix B.
The ${\cal MT}$ model can be obtained as a special limit case of the proposed $\mathcal{H}_{\kappa,\mu,a,d}$ model after two suitable compact support reparameterizations. To obtain the first parameterization, the following theorem provides sufficient conditions of equivalence between two Gaussian measures with covariances $\sigma_{0}^{2}{\cal MT}_{\nu,\alpha}$ and $\sigma_{1}^{2}\mathcal{H}_{\kappa,\mu,a,d}$, respectively. The proof can be found in Appendix A.
@@
Let $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})$ and $P(\sigma_{1}^{2}\mathcal{H}_{\kappa,\mu,a,d})$ be two zero-mean Gaussian measures. If $\nu=\kappa+\frac{1}{2}$ with $\nu>-\frac{1}{2}$ and
@@
then for any bounded infinite set $D\subset\mathbb{R}^{d}$, $d=1,2,3$, $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})\equiv P(\sigma_{1}^{2}\mathcal{H}_{\kappa,\mu,a,d})$ on the paths of $Z(\boldsymbol{s}),\boldsymbol{s}\in D$, where
$B(\kappa,\mu,d)=\frac{2^{2\kappa+1}\Gamma(\frac{\mu+1}{2}+\kappa)\Gamma(\frac{\mu+d+1}{2}+2\kappa)}{\Gamma(\frac{\mu}{2})\Gamma(\frac{\mu+d}{2}+\kappa)}.$ (24)
Note that the equivalence condition (23) sharing a common variance parameter and taking into account the relation between the smoothness parameters $\nu=\kappa+\frac{1}{2}$, can be written as
@@
If we use the previous equation as a compact support reparameterization in the ${\cal H}$ model, then, using the same arguments as in Bevilacqua et al., (2022), it can be shown that
@@
with uniform convergence on $[0,\infty)$. In addition, following the same arguments in Emery et al., (2025), it can be shown that
$\lim_{\mu\to\infty}{\cal H}_{\kappa,\mu,\mu\alpha,d}(x)={\cal MT}_{\kappa+\frac{1}{2},\alpha}(x),\quad\kappa\geq-\frac{1}{2},$ (26)
with uniform convergence on $[0,\infty)$. As a consequence, the two reparameterized versions of the ${\cal H}$ model can be viewed as generalizations of the ${\cal MT}$ model with an additional parameter $\mu$ that, for given smoothness and spatial dependence parameters, allows to obtain compactly ($\mu<\infty$) or globally ($\mu=\infty$) supported covariance functions. Thus, under this parameterization, $\mu$ can be fixed by the user, particularly when dealing with large datasets, to achieve sparse covariance matrices and obtain computational benefits.
An illustration of (25) can be found in Figure 4 where the ${\cal H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ model with increasing values of $\mu$ is compared with the ${\cal MT}_{\kappa+\frac{1}{2},\alpha}$ model when $d=2$.
Refer to caption Refer to caption
Figure 4: The ${\cal H}_{\kappa,\mu,b(\alpha,\mu,\kappa,2),2}$ for $\mu=1,2,10$ when $\kappa=0$ (left) and $\kappa=1$ (right). The red line is the ${\cal MT}_{\kappa+\frac{1}{2},\alpha}$ correlation model ($\alpha=0.2$ in both cases).
It should be emphasized that, under this compact support reparameterization, the role of the shape parameter $\mu$ changes significantly. Let us assume, without loss of generality, that the scale dependence parameter $\alpha$ is equal to $1$. Under this assumption, the parameter $\mu$ (together with the smoothness parameter and the dimension $d$) determines the compact support. For large values of $\mu$, the compact support becomes larger, and both the ${\cal H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ and ${\cal H}_{\kappa,\mu,\mu,d}$ models approach the ${\cal MT}_{\kappa+\frac{1}{2},1}$ model. In this case, the maximum integral range is the one associated with the ${\cal MT}$ model in Equation (17).
Conversely, as $\mu$ approaches its lower bound, the compact support shrinks, and a potential smaller compact support results in a more flexible model. If we compare the compact support of $\mathcal{H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ with that of $\mathcal{GW}_{\kappa,\mu,g(1,\mu,\kappa)}$ (see Equation 9) and given that $\mu$ is lower bounded by $1$ for the $\mathcal{H}$ model and by $\mu_{\min}(\kappa,d)$ for the $\mathcal{GW}$ model, it can be shown that
$b(1,1,\kappa,d)\leq g(1,\mu_{\min}(\kappa,d),\kappa)$
when $d\geq 2$ or $\kappa\geq 0$ and $b(1,1,\kappa,d)>g(1,\mu_{\min}(\kappa,d),\kappa)$ when $d=1$ and $-\frac{1}{2}<\kappa<0$. This implies that the $\mathcal{H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ model achieves a smaller compact support compared to the $\mathcal{GW}_{\kappa,\mu,g(1,\mu,\kappa)}$ model for a significant portion of the parameter space. As a consequence, the $\mathcal{H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ model is more flexible from this perspective.
On the other hand, if we compare the smaller compact supports of $\mathcal{H}_{\kappa,\mu,\mu,d}$ and $\mathcal{GW}_{\kappa,\mu,\mu}$, then we reach the same conclusions since $1\leq\mu_{\min}(\kappa,d)$ when $d\geq 2$ or $\kappa\geq 0$ and $1>\mu_{\min}(\kappa,d)$ when $d=1$ and $-\frac{1}{2}<\kappa<0$.
We also compare the smaller compact supports that can be achieved by the ${\cal H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ and ${\cal H}_{\kappa,\mu,\mu,d}$ models, which are $b(1,1,\kappa,d)$ and $1$, respectively. Figure 5 illustrates the compact support $b(1,1,\kappa,d)$ for $-0.5<\kappa<1.5$ and $d=1,2,3,4,5$, compared to the reference value $1$. It can be observed that only for a small region of the smoothness parameter space and $d=1,2,3,4$, we obtain $b(1,1,\kappa,d)<1$.
Figure 5: The smaller compact support $b(1,1,\kappa,d)$ that can be achieved by the model ${\cal H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ for $d=1,2,3,4,5$ and $-0.5<\kappa<1.5$. The red line is the smaller compact support that can be achieved by the ${\cal H}_{\kappa,\mu,\mu,d}$ model.
Summarizing our results, when comparing the two reparameterized versions of the $\mathcal{H}$ model with the analogous versions of the $\mathcal{GW}$ model that share the same smoothness parameter $\kappa$ and converge to the same $\mathcal{MT}$ model as the shape parameter $\mu$ becomes very large, it turns out that the reparameterized $\mathcal{H}$ models can achieve smaller compact supports. In addition the ${\cal H}_{\kappa,\mu,\mu,d}$ model can achieve smaller compact supports with respect to the ${\cal H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ model.
@@ 6 Asymptotic properties of the maximum likelihood estimator of the Hypergeometric model
When considering the asymptotic properties of the ML estimator for the parameters of the $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ model, two different scenarios can be examined. Under increasing domain asymptotics, the parameters $(\sigma^{2},\kappa,\mu,a)$ can be estimated consistently. Furthermore, the asymptotic distribution of the ML estimator is Gaussian, with a covariance matrix equal to the inverse of the Fisher information matrix (Mardia and Marshall,, 1984).
Under fixed-domain, using the results of Theorem 1 and taking into account the parameterization (15) with $l=\frac{d}{2}+\kappa$, the equivalence condition for the two Gaussian measures $P(\sigma_{i}^{2}{\cal H}_{\kappa,\mu_{i},a_{i},d})$, $i=0,1$, sharing a common smoothness parameter, when $d=1,2,3$ and provided that $\min(\mu_{0},\mu_{1})\geq 1+\frac{d}{2}$, reads:
$\frac{\sigma_{0}^{2}B(\kappa,\mu_{0},d)}{a_{0}^{2\kappa+1}}=\frac{\sigma_{1}^{2}B(\kappa,\mu_{1},d)}{a_{1}^{2\kappa+1}},$ (27)
As a consequence only the microergodic parameter $\sigma^{2}B(\kappa,\mu,d)/a^{2\kappa+1}$ can be estimated consistently. On the other hand, assuming a fixed known shape parameter $\mu$, the microergodic parameter is given by $\sigma^{2}/a^{2\kappa+1}$.
In what follows, we provide the asymptotic properties of ML estimator of the microergodic parameter $\sigma^{2}/a^{2\kappa+1}$. In particular, we will show that the microergodic parameter can be estimated consistently, and then we will assess the asymptotic distribution of the ML estimator.
Let $D$ be a bounded subset of $\mathbb{R}^{d}$ and let $\boldsymbol{Z}_{n}=(Z(\boldsymbol{s}_{1}),\ldots,Z(\boldsymbol{s}_{n}))^{\prime}$ be a finite realization of $Z(\boldsymbol{s})$, $\boldsymbol{s}\in D$, a zero-mean stationary Gaussian RF with a given parametric covariance function $\sigma^{2}\phi(\|\cdot\|;\boldsymbol{\tau})$, with $\sigma^{2}>0$, $\boldsymbol{\tau}$ a parameter vector and $\phi$ a member of the class $\Phi_{d}$, with $\phi(0;\boldsymbol{\tau})=1$.
We then write $R_{n}(\boldsymbol{\tau})=[\phi(\|\boldsymbol{s}_{i}-\boldsymbol{s}_{j}\|;\boldsymbol{\tau})]_{i,j=1}^{n}$ for the associated correlation matrix. The Gaussian log-likelihood function is defined as:
$\mathcal{L}_{n}(\sigma^{2},\boldsymbol{\tau})=-\frac{1}{2}\left(n\log(2\pi\sigma^{2})+\log(|R_{n}(\boldsymbol{\tau})|)+\frac{1}{\sigma^{2}}\boldsymbol{Z}_{n}^{\prime}R_{n}(\boldsymbol{\tau})^{-1}\boldsymbol{Z}_{n}\right).$ (28)
Under the $\cal H$ model, the Gaussian log-likelihood is obtained with $\phi(\cdot;\boldsymbol{\tau})\equiv\mathcal{H}_{\kappa,\mu,a,d}$ with $\boldsymbol{\tau}=(\kappa,\mu,a)^{\prime}$. Since in what follows $\mu$ and $\kappa$ are assumed known and fixed, for notation convenience, we write $\tau=a$.
Let $\hat{\sigma}^{2}_{n}$ and $\hat{a}_{n}$ be the ML estimator obtained maximizing $\mathcal{L}_{n}(\sigma^{2},a)$ for a fixed $\mu$ and $\kappa$. We consider two types of estimators. The first maximizes (28) with respect to $\sigma^{2}$ for a fixed arbitrary compact support $a>0$, obtaining the following estimator
$\hat{\sigma}_{n}^{2}(a)=\operatorname*{arg\,max}_{\sigma^{2}}\mathcal{L}_{n}(\sigma^{2},a)=\boldsymbol{Z}_{n}^{\prime}R_{n}(a)^{-1}\boldsymbol{Z}_{n}/n.$ (29)
Here $R_{n}(a)$ is the correlation matrix coming from the $\mathcal{H}_{\kappa,\mu,a,d}$ model. The following result establishes strong consistency and asymptotic distribution of the ML estimator $\hat{\sigma}_{n}^{2}(a)/a^{(1+2\kappa)}$ of the microergodic parameter. We omit the proof, as it follows the same steps as Theorem 8 in Bevilacqua et al., (2019).
@@
Let $Z(\boldsymbol{s})$, $\boldsymbol{s}\in D\subset\mathbb{R}^{d}$, $d=1,2,3$, be a zero-mean Gaussian RF with $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ covariance model with $\mu\geq 1+d/2$. Suppose $(\sigma_{0}^{2},a_{0})\in(0,\infty)\times(0,\infty)$. For a fixed $a>0$, let $\hat{\sigma}_{n}^{2}(a)$ as defined in Equation (29). Then, as $n\to\infty$,
@@
    $\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}\stackrel{{\scriptstyle a.s}}{{\longrightarrow}}\sigma_{0}^{2}(a_{0})/a_{0}^{1+2\kappa}$ and
@@
    $\sqrt{n}(\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}-\sigma_{0}^{2}(a_{0})/a_{0}^{1+2\kappa})\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}\mathcal{N}(0,2(\sigma_{0}^{2}(a_{0})/a_{0}^{1+2\kappa})^{2})$.
The second type of estimation technique considers the joint maximization of (28) with respect to $(\sigma^{2},a)\in(0,\infty)\times I$ where $I=[a_{L},a_{U}]$ and $0<a_{L}<a_{U}<\infty$. The solution of this optimization problem is given by $(\hat{\sigma}_{n}^{2}(\hat{a}_{n}),\hat{a}_{n})$ where
$\hat{\sigma}_{n}^{2}(\hat{a}_{n})=\boldsymbol{Z}_{n}^{\prime}R_{n}(\hat{a}_{n})^{-1}\boldsymbol{Z}_{n}/n$
and $\hat{a}_{n}=\operatorname*{arg\,max}_{a\in I}\mathcal{PL}_{n}(a)$. Here $\mathcal{PL}_{n}(a)$ is the profile log-likelihood:
$\mathcal{PL}_{n}(a)=-\frac{1}{2}\left(\log(2\pi)+n\log(\hat{\sigma}_{n}^{2}(a))+\log|R_{n}(a)|+n\right).$ (30)
In order to establish strong consistency and asymptotic distribution of the sequence of random variables $\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}$, we use the following lemma that establishes the monotone behavior of $\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}$ when viewed as a function of $a\in I$ under specific conditions on the $\mu$ parameter. The proof can be found in Appendix A.
@@
Let $S_{n}=\{\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{n}\in D\subset\mathbb{R}^{d}\}$ denote any set of distinct locations. For any $a_{1}<a_{2}\in I$ and for each $n$, $\hat{\sigma}_{n}^{2}(a_{1})/a_{1}^{1+2\kappa}\leq\hat{\sigma}_{n}^{2}(a_{2})/a_{2}^{1+2\kappa}$ if and only if $\mu\geq 4$.
Finally, the following theorem establishes the asymptotic properties of the ML estimator $\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}$. We omit the proof, as it follows the same steps as Theorem 9 in Bevilacqua et al., (2019) using the results in Lemma 1.
@@
Let $Z(\boldsymbol{s})$, $\boldsymbol{s}\in D\subset\mathbb{R}^{d}$, $d=1,2,3$, be a zero-mean Gaussian RF with $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ covariance model with $\mu\geq 4$. Suppose $(\sigma_{0}^{2},a_{0})\in(0,\infty)\times I$ where $I=[a_{L},a_{U}]$ with $0<a_{L}<a_{U}<\infty$. Let $(\hat{\sigma}_{n}^{2},\hat{a}_{n})^{\prime}$ maximize (28) over $(0,\infty)\times I$. Then as $n\to\infty$,
@@
    $\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}\stackrel{{\scriptstyle a.s}}{{\longrightarrow}}\sigma_{0}^{2}(a_{0})/a_{0}^{1+2\kappa}$ and

### v2

@@ 5 Some special cases and bridges to the Matérn model
The computation of the correlation function $\mathcal{H}_{\kappa,\mu,a,d}$ requires the evaluation of the Gauss hypergeometric function. Although efficient numerical implementations are available (Johansson, 2019), this may be less familiar to spatial statisticians and machine learning practitioners accustomed to closed-form parametric correlation models.
As in the $\mathcal{MT}$ and $\mathcal{GW}$ families, substantial simplifications arise in certain special cases. The following theorem considers the case $\kappa=k\in\mathbb{N}_{0}$, $d=1$, and $\mu\geq 1$.
@@
If $\kappa=k$ is a nonnegative integer, $d=1$, and $\mu\geq 1$, then
$\begin{split}\mathcal{H}_{k,\mu,a,1}(x)&=\frac{\sqrt{\pi}\,\Gamma\!\left(\frac{\mu}{2}+2k+1\right)}{\Gamma\!\left(k+\tfrac{1}{2}\right)}\left(1-\frac{x}{a}\right)^{\mu+2k}\left(\frac{x}{a}\right)^{k}\\ &\quad\times\sum_{n=0}^{k}\frac{(n+k)!}{2^{2n}n!\,\Gamma\!\left(n+\frac{\mu}{2}+k+1\right)(k-n)!}\left(\frac{(1-\frac{x}{a})^{2}}{\frac{x}{a}}\right)^{n}.\end{split}$
General closed forms for arbitrary $d>1$, $\mu\geq 1$, and integer $\kappa$ are more difficult to obtain. Nevertheless, additional simplifications can be derived using known identities of the Gauss hypergeometric function. For instance, when $d$ is odd, $\mu\geq 1$, and $\kappa=0$, the following closed-form expression holds.
@@
Let $d$ be an odd integer and $\mu\geq 1$. If $\kappa=0$, then
$\mathcal{H}_{0,\mu,a,d}(x)=C\left(1-\frac{x}{a}\right)_{+}^{\mu+\frac{d-1}{2}}\sum_{n=0}^{\frac{d-1}{2}}\frac{\left(\frac{1-d}{2}\right)_{n}\left(\frac{d+1}{2}\right)_{n}\left(1-\frac{x}{a}\right)_{+}^{n}}{2^{n}n!\,\Gamma\!\left(n+\frac{d-1}{2}+\mu+1\right)},$
where
$C=\frac{2^{\mu+\frac{d-1}{2}}\Gamma\!\left(\frac{\mu+1}{2}\right)\Gamma\!\left(\frac{\mu+d+1}{2}\right)}{\sqrt{\pi}}.$
Additional closed-form expressions, including spherical models, are reported in Appendix B.
The Matérn model can be recovered as a limiting case of $\mathcal{H}_{\kappa,\mu,a,d}$ under an appropriate compact-support reparameterization. To establish this result, we first characterize the equivalence between the Gaussian measures $P(\sigma_{0}^{2}\mathcal{MT}_{\nu,\alpha})$ and $P(\sigma_{1}^{2}\mathcal{H}_{\kappa,\mu,a,d})$. The proof is provided in Appendix A.
@@
Let $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})$ and $P(\sigma_{1}^{2}{\cal H}_{\kappa,\mu,a,d})$ be two zero-mean Gaussian measures, and let $\kappa>-\frac{1}{2}$ and $\nu=\kappa+\frac{1}{2}$. For any bounded infinite set $D\subset\mathbb{R}^{d}$, $d=1,2,3$, the Gaussian measures $P(\sigma_{0}^{2}{\cal MT}_{\nu,\alpha})$ and $P(\sigma_{1}^{2}{\cal H}_{\kappa,\mu,a,d})$ are equivalent on the paths of $Z(\mathbf{s})$, $\mathbf{s}\in D$, if and only if
@@
where
$B(\kappa,\mu,d)=\frac{2^{2\kappa+1}\Gamma\!\left(\frac{\mu+1}{2}+\kappa\right)\Gamma\!\left(\frac{\mu+d+1}{2}+2\kappa\right)}{\Gamma\!\left(\frac{\mu}{2}\right)\Gamma\!\left(\frac{\mu+d}{2}+\kappa\right)}.$ (20)
Assuming a common variance and imposing the relation $\nu=\kappa+\frac{1}{2}$, the equivalence condition (19) is equivalently expressed as
@@
Using this compact-support reparameterization and following the arguments in Bevilacqua et al. (2022), one obtains
@@
with uniform convergence on $0,\infty)$.
[…]
It is worth emphasizing that, under this compact-support reparameterization, the role of the shape parameter $\mu$ changes substantially. For fixed smoothness $\kappa$ and scale parameter $\alpha=1$, $\mu$ determines the compact support.
As $\mu\to\infty$, ${\cal H}_{\kappa,\mu,b(1,\mu,\kappa,d),d}$ converges to the Matérn model ${\cal MT}_{\kappa+\frac{1}{2},1}$, with maximal integral range given by (13). Conversely, as $\mu$ approaches its lower bound, the compact support decreases, potentially yielding more flexible local dependence structures.
A comparison with the corresponding $\mathcal{GW}_{\kappa,\mu,g(1,\mu,\kappa)}$ model (see (6)) shows that, over a substantial portion of the parameter space, the reparameterized $\mathcal{H}$ models can achieve smaller compact supports. In particular, when $d\geq 2$ or $\kappa\geq 0$,
$b(1,1,\kappa,d)\leq g(1,\mu_{\min}(\kappa,d),\kappa),$
where $\mu_{\min}(\kappa,d)=\kappa+\frac{d+1}{2}$. The case $d=1$ and $\frac{1}{2}<\kappa<0$ is excluded here, as the exact lower bound $\mu_{\min}$ for the $\mathcal{GW}$ model remains an open problem in that regime.
@@ 6 Asymptotic properties of the maximum likelihood estimator of the Hypergeometric model
We study the asymptotic properties of the maximum likelihood (ML) estimator for the parameters of the model $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ under both increasing-domain and fixed-domain asymptotics.
Under increasing-domain asymptotics, the parameters $(\sigma^{2},\kappa,\mu,a)$ are consistently estimable, and the ML estimator is asymptotically Gaussian with covariance matrix equal to the inverse Fisher information (Mardia and Marshall, 1984).
Under fixed-domain asymptotics, the situation differs. Using Theorem 1 and the parameterization (11) with $l=\frac{d}{2}+\kappa$, the equivalence condition for two Gaussian measures $P(\sigma_{i}^{2}{\cal H}_{\kappa,\mu_{i},a_{i},d})$, $i=0,1$, sharing the same smoothness parameter, for $d=1,2,3$ and $\min(\mu_{0},\mu_{1})\geq 1+\frac{d}{2}$, is
$\frac{\sigma_{0}^{2}B(\kappa,\mu_{0},d)}{a_{0}^{2\kappa+1}}=\frac{\sigma_{1}^{2}B(\kappa,\mu_{1},d)}{a_{1}^{2\kappa+1}}.$ (22)
Hence only the microergodic parameter
$\frac{\sigma^{2}B(\kappa,\mu,d)}{a^{2\kappa+1}}$
is consistently estimable. If $\mu$ is fixed and known, this reduces to $\sigma^{2}/a^{2\kappa+1}$.
Let $D\subset\mathbb{R}^{d}$ be bounded and $\mathbf{Z}_{n}=(Z(\mathbf{s}_{1}),\ldots,Z(\mathbf{s}_{n}))^{\prime}$ a realization of a zero-mean stationary Gaussian RF with covariance $\sigma^{2}\phi(\|\cdot\|;\boldsymbol{\tau})$, $\phi\in\Phi_{d}$, $\phi(0;\boldsymbol{\tau})=1$. Denote by
$R_{n}(\boldsymbol{\tau})=[\phi(\|\mathbf{s}_{i}-\mathbf{s}_{j}\|;\boldsymbol{\tau})]_{i,j=1}^{n}$
the correlation matrix. The Gaussian log-likelihood is
$\mathcal{L}_{n}(\sigma^{2},\boldsymbol{\tau})=-\frac{1}{2}\Big(n\log(2\pi\sigma^{2})+\log|R_{n}(\boldsymbol{\tau})|+\frac{1}{\sigma^{2}}\mathbf{Z}_{n}^{\prime}R_{n}(\boldsymbol{\tau})^{-1}\mathbf{Z}_{n}\Big).$ (23)
Under the $\mathcal{H}$ model, $\phi(\cdot;\boldsymbol{\tau})=\mathcal{H}_{\kappa,\mu,a,d}$ with $\boldsymbol{\tau}=(\kappa,\mu,a)^{\prime}$. In what follows, $\kappa$ and $\mu$ are assumed fixed and known, so we write $\tau=a$.
Let $\hat{\sigma}_{n}^{2}(a)$ denote the ML estimator of $\sigma^{2}$ for fixed $a>0$:
$\hat{\sigma}_{n}^{2}(a)=\mathbf{Z}_{n}^{\prime}R_{n}(a)^{-1}\mathbf{Z}_{n}/n.$ (24)
The following result establishes strong consistency and asymptotic normality of $\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}$. The proof follows the same arguments as Theorem 8 in Bevilacqua et al. (2019).
@@
Let $Z(\mathbf{s})$, $\mathbf{s}\in D\subset\mathbb{R}^{d}$, $d=1,2,3$, be a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$, $\mu\geq 1+\frac{d}{2}$. Suppose $(\sigma_{0}^{2},a_{0})\in(0,\infty)^{2}$. Then, for fixed $a>0$, as $n\to\infty$,
@@
    $\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\sigma_{0}^{2}/a_{0}^{1+2\kappa}$;
@@
    $\sqrt{n}\big(\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}-\sigma_{0}^{2}/a_{0}^{1+2\kappa}\big)\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}\mathcal{N}\big(0,2(\sigma_{0}^{2}/a_{0}^{1+2\kappa})^{2}\big)$.
We next consider joint maximization of (23) with respect to $(\sigma^{2},a)\in(0,\infty)\times I$, $I=[a_{L},a_{U}]$, $0<a_{L}<a_{U}<\infty$. Let
$\hat{\sigma}_{n}^{2}(\hat{a}_{n})=\mathbf{Z}_{n}^{\prime}R_{n}(\hat{a}_{n})^{-1}\mathbf{Z}_{n}/n,\quad\hat{a}_{n}=\arg\max_{a\in I}\mathcal{PL}_{n}(a),$
where $\mathcal{PL}_{n}(a)$ is the profile log-likelihood
$\mathcal{PL}_{n}(a)=-\frac{1}{2}\big(\log(2\pi)+n\log\hat{\sigma}_{n}^{2}(a)+\log|R_{n}(a)|+n\big).$ (25)
The following lemma establishes a monotonicity property of $\hat{\sigma}_{n}^{2}(a)/a^{1+2\kappa}$ as a function of $a$. Its proof is given in Appendix A.
@@
Let $S_{n}=\{\mathbf{s}_{1},\ldots,\mathbf{s}_{n}\}\subset D$. For any $a_{1}<a_{2}$ in $I$,
$\frac{\hat{\sigma}_{n}^{2}(a_{1})}{a_{1}^{1+2\kappa}}\leq\frac{\hat{\sigma}_{n}^{2}(a_{2})}{a_{2}^{1+2\kappa}}\quad\text{if and only if}\quad\mu\geq 4.$
The next theorem establishes the asymptotic properties of $\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}$. The proof follows Theorem 9 in Bevilacqua et al. (2019) combined with Lemma 1.
@@
Let $Z(\mathbf{s})$, $\mathbf{s}\in D\subset\mathbb{R}^{d}$, $d=1,2,3$, be a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$, $\mu\geq 4$. Suppose $(\sigma_{0}^{2},a_{0})\in(0,\infty)\times I$. Then, as $n\to\infty$,
@@
    $\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\sigma_{0}^{2}/a_{0}^{1+2\kappa}$;